OpenStudy (anonymous):

All pairs integers (m, n) such that 2^(m+1) + 3^(n+1) is perfect square

OpenStudy (zzr0ck3r):

what is the question?

OpenStudy (anonymous):

Find all order pairs (m,n) which satisfy that eq

OpenStudy (anonymous):

2^(m+1) + 3^(n+1) = k^2

OpenStudy (anonymous):

By using trials i got one point (3,1). But i dont know with other solutions

OpenStudy (anonymous):

What about you @zzr0ck3r

OpenStudy (rational):

*

OpenStudy (anonymous):

is there a nice way, @rational :)

OpenStudy (rational):

not getting ideas sry @dan815 @Kainui

OpenStudy (rational):

here is my attempt assuming \(m,n\) are positive : \[2^{m+1}+3^{n+1}=k^2\] Notice that \(2^{m+1}\equiv 0\pmod{4}\) and \(3^{n+1}\equiv (-1)^{n+1}\pmod{4}\). Recall the fact that any perfect square is congruent to 0,1 in mod 4, therefore \(n+1\) must be even. let \(n+1=2r\) : \[2^{m+1}+3^{2r}=k^2\] rearrange the equation and get \[2^{m+1} = k^2-3^{2r}\] factoring the right hand side \[2^{m+1} = (k+3^r)(k-3^r)\]

OpenStudy (kainui):

I was trying but not able to get it to come together by relating the fact that every square is the sum of consecutive odd integers and that two consecutive triangular numbers also make a square with the binomial theorem and it written in a weird kind of way: \[\left( \frac{5+1}{2}\right)^{n+1} + \left( \frac{5-1}{2}\right)^{m+1} = k^2\] I'll have to come back and think about this, interesting though!